# Accurate BGV Parameters Selection: Accounting for Secret and Public Key Dependencies in Average-Case Analysis

### v2

@@ Abstract
In this paper, we provide a new method to estimate noise growth, closely aligning with experimental results and forming the basis for parameter selection that ensures correctness and improves efficiency.
@@ 1 Introduction
The homomorphic encryption schemes currently in use base their security on the presumed intractability of the Learning with Errors (LWE) problem, [26], and its ring variant (RLWE) [23]. Informally, the decisional version of RLWE consists of distinguishing polynomial equations $(a,b=s\cdot a+e)\in\mathcal{R}_{q}\times\mathcal{R}_{q}$, perturbed by small noise $e$ (also called error), from uniform random tuples from $\mathcal{R}_{q}\times\mathcal{R}_{q}$, where $\mathcal{R}_{q}=\mathbb{Z}_{q}[x]/\langle x^{n}+1\rangle$ and $q$ is a positive integer.
@@
Interest in this method, initially applied in the TFHE scheme [9], and subsequently in the CKKS [12], BGV [14, 25] and BFV [2] schemes, grew due to a recognized discrepancy between the estimates based on worst-case technique and experimental data, as highlighted in [13]. The introduction of the average-case approach, as seen in [2, 14], offers a potential resolution to these disparities, indeed, with this method, it is possible to compute a tight probabilistic upper bound.
However, the heuristics used for the BGV [25] and CKKS [12] schemes often underestimate the noise growth due to the assumption of the noises independence, leading to imprecise bounds, as also pointed out in [1, 12, 25]. Such underestimates lead to two potential issues: first, the ciphertext is not correctly decrypted with non-negligible probability due to excessive noise and, second, the scheme is exposed to security vulnerabilities, as shown in recent papers [5, 6].
@@
In this paper, we propose the first average-case noise analysis for BGV that does not provide underestimates, taking in account the dependencies introduced by the common secret and public key. We extend the approach of BFV [2], where the authors consider the fact that the errors are not independent among each other, and introduce a function $F$ to “correct” the product of the variances for homomorphic multiplication. We notice the additional dependencies given by the public key and employ a second correcting function to address its contribution. The results obtained in this study suggest that this approach leads to significant improvements in noise analysis.
@@
To account for these dependencies, we applied the correction function $F$ introduced in [2]. It is important to note that the case of BGV is more complex than that of BFV. In BFV, the dependency due to $e$ appears only in a negligible term. In contrast, for BGV, we must take into account the dependencies on both $s$ and $e$. This requires the use of two distinct correction functions to accurately model and mitigate these dependencies.
The paper is structured as follows. Section 2 introduces essential definitions and fundamental properties that are instrumental for understanding both our contribution and the scheme more broadly. Section 3 provides a concise overview of the main features and structure of the BGV scheme. In Section 4, we present our key results concerning the behavior of the error term and its growth under homomorphic operations. Section 5 then demonstrates how the findings from Section 4 can be leveraged to estimate error growth in fixed-operation circuits, laying the groundwork for novel approaches to selecting the scheme’s initial parameters. Finally, Section 6 concludes the paper and outlines possible directions for future research inspired by our results.
@@ 2 Preliminaries
In this section, we define the general notations that we will use in the remainder of the work.
In our analysis we will describe the leveled version of the BGV scheme, noting that this implies that the parameters of the scheme depend (polynomially) on the depth of the circuits that the scheme is capable of evaluating [4]. It should be noted that this version is also the most widely adopted because it does not require the usage of the bootstrapping technique, which is highly complex from a computational perspective. Before recalling the BGV scheme, some preliminary parameters must be fixed.
Let $L$ denote the number of levels of the arithmetic circuit that the scheme must be able to evaluate. Moreover, assume that $m$ is a suitably chosen integer. The notation $\mathcal{R}_{d}$, for a fixed $d\in\mathbb{Z}$, will denote the ring $\mathcal{R}_{d}=\mathbb{Z}_{d}[x]/\langle\Phi_{m}(x)\rangle$, where $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial, a definition of which is briefly recalled below.
Definition 1
Given a positive integer $m$, the $m$-th cyclotomic polynomial is defined as
$\Phi_{m}(x)=\prod_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}(x-\zeta^{j}),$
where $\zeta$ is a primitive $m$-th root of unity. This polynomial has degree $\phi(m)$, where $\phi$ denotes the Euler function.
Eventually, we will denote with $\mathcal{R}$ the ring $\mathcal{R}=\mathbb{Z}[x]/\langle\Phi_{m}(x)\rangle$. To define the plaintext and ciphertext spaces, $L+1$ moduli are selected as follows:
  • An integer $t$, denoted as the plaintext modulus, is chosen such that $t\equiv 1\mod m$.
  • $L$ integers $q_{l}$, for $l\in\{0,\dots,L-1\}$, are defined as:
    $q_{l}=\prod_{j=0}^{l}p_{j},$
    where the $p_{j}$ are primes satisfying $p_{j}\equiv 1\mod m$ for $j=0,\dots,L-1$ and $p_{1},...,p_{L-2}$ have approximately the same size. Each $q_{L-1-l}$ is referred to as the ciphertext modulus for the $l$-th level.
According to these notations, we can define the plaintext and ciphertext spaces, respectively, as $P=\mathcal{R}_{t}$, $C=\mathcal{R}_{q_{l}}\times R_{q_{l}}$.
Due to the ciphertext coefficients being reduced according to different moduli at each level, BGV is commonly referred to as scale-dependent.
@@
We use lowercase letters such as $x$ to denote polynomials, and bold symbols like $\bm{x}$ to represent vectors of polynomials. Given a polynomial $x\in\mathcal{R}_{n}$, we use the notation $[x]_{n}$ to refer to the centered representation of $x$ modulo $n$ i.e., the representative of $x$ with coefficients in $[-\frac{n}{2},\frac{n}{2})$. In general, unless otherwise specified, we always assume the coefficients of polynomials in $\mathcal{R}_{d}$ to be centered with respect to the modulus $d$.
Moreover, given a random polynomial $p\in\mathcal{R}$ and a probabilistic distribution $X$, the notation $p\leftarrow X$ is used to indicate that each coefficient of $p$ is chosen independently according to $X$.
Some distributions that will be frequently considered are the following:
  • $\mathcal{U}_{q}$ as the uniform distribution over $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$;
  • $\mathcal{N}(0,\sigma^{2})$ as the normal distribution, also referred to as the Gaussian distribution, over $\mathbb{R}$, with mean 00 and variance $\sigma^{2}$;
  • $\mathcal{DG}_{q}(\sigma^{2})$ as the discrete Gaussian distribution, which involves sampling a value according to $\mathcal{N}(0,\sigma^{2})$, rounding it to the nearest integer and then reducing it modulo $q$.
    Moreover, the representative modulo $q$ is taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$;
For the BGV scheme the notation $\chi_{e},\chi_{s}$ will be adopted in order to indicate the distribution of the error for a RLWE instance and the secret key coefficients, respectively. Typically, $\chi_{e}$ is a discrete Gaussian distribution with a suitable standard deviation, while for the secret key, the ternary uniform distribution $\mathcal{U}_{3}$ is usually considered. However, in some special cases, where bootstrapping is needed, the choice for the secret key distribution falls on the Hamming Weight distribution, a definition of which is recalled below [2].
Definition 2
The Hamming Weight distribution $\mathcal{HWT}_{n,h}$ consists of selecting uniformly at random a vector of $n$ entries in $\{-1,0,1\}$, where exactly $h<n$ entries are non-zero.
To ensure the correctness of the scheme, it is important to underline that the error added to the plaintext during encryption must be sufficiently small [4].
This will become more apparent through the presentation of the BGV functions.
@@
For the purposes of our subsequent analysis, it is helpful to recall that for two polynomials $r,s\in\mathcal{R}$, and denoting by $r|_{i}$ and $s|_{i}$ their coefficients with respect to $x^{i}$, the $i$-th coefficient of their product is given by [2]
$rs|_{i}=\sum_{j=0}^{\phi(m)-1}\xi(i,j)r|_{j}s|_{i-j},$ (1)
where
$\xi(i,j)=\left\{\begin{matrix}1&\text{\hskip 51.21504ptfor }i-j\in[0,\phi(m))\\[5.0pt] -1&\text{otherwise}\end{matrix}\right.$
Furthermore, the following notation will be adopted.
Notation 1
Given a polynomial $p\in\mathcal{R}_{q}$, for a fixed integer $q$, the variance of its coefficients, namely $\mathsf{Var}(p|_{i})$, is denoted as $V_{p}$.
Proposition 1
Let $p,q$ be two independent polynomials in $\mathcal{R}_{q}$, where $q$ is a fixed integer, and let $k$ be an integer in $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$. Moreover, assume that the coefficients of each polynomial are independent and identically distributed, with 00-mean. Then,
  • a.
    $V_{p+q}=V_{p}+V_{q}$
  • b.
    $V_{kp}=k^{2}V_{p}$
  • c.
    $V_{p\cdot q}=\phi(m)V_{p}\cdot V_{q}$
@@
Finally, the variance values for some common distributions, which will often be employed in the BGV scheme, are:
@@ 3 The BGV scheme
Once the parameters have been specified, the framework of the BGV scheme can be defined according to three main functions: Key Generation, Encryption and Decryption.
3.1 Key Generation
The key generation function is responsible for producing:
  • the secret key, represented by a polynomial in $\mathcal{R}_{q_{L-1}}$;
  • the public key, which consists of a pair of polynomials, each belonging to $\mathcal{R}_{q_{L-1}}$;
To achieve this, three polynomials in $\mathcal{R}_{q_{L-1}}$ are generated according to the following distributions: $s\leftarrow\chi_{s}$ , $a\leftarrow\mathcal{U}_{q_{L-1}}$ and $e\leftarrow\chi_{e}$. The resulting keys are given by
$\begin{cases}sk&=s\\ pk&=(b,a)\equiv(-a\cdot s+te,a)\mod q_{L-1}\end{cases}$
3.2 Encryption
@@
Given the plaintext $m\in\mathcal{R}_{t}$ and the public key $pk=(b,a)$, the encryption function returns as output $(\bm{c},q_{l},\nu_{clean})$ where:
  • $\bm{c}\in\mathcal{R}_{q_{l}}\times\mathcal{R}_{q_{l}}$ is the actual ciphertext defined as
    $\bm{c}=(c_{0},c_{1})\equiv(b\cdot u+te_{0}+m,a\cdot u+te_{1})\mod q_{l},$
    where $u,e_{0},e_{1}\in\mathcal{R}_{q_{l}}$, with coefficients distributed as $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
  • $q_{l}$ represents the ciphertext modulus for the level at which computations are performed. Specifically, as the encryption function is computed in the first level 00, the corresponding $q_{l}$ is $q_{L-1}$.
  • $\nu$ is the critical quantity associated to the ciphertext and is defined as
    $\nu=[c_{0}+c_{1}\cdot s]_{q_{l}}.$
We refer to the triple $\tilde{c}=(\bm{c},q_{l},\nu_{clean})$ with the term extended ciphertext.
Broadly speaking, the encryption of the plaintext message results in a pair of ciphertext values: the first encodes the masked plaintext using the public key, while the second supplies auxiliary information necessary for the receiver to recover the original message, provided they possess the secret key.
Regarding the critical quantity, it is important to emphasize that this value is strictly related to the error introduced by the encryption operator. Thus, it essentially determines whether the ciphertext can be correctly decrypted or not. For this reason, its behavior is analyzed to gain a clear understanding of the functioning of the scheme.
3.3 Decryption
At any level, it should be possible to recover the original plaintext using the decryption function.
@@ 3.3 Decryption
Given the extended ciphertext $(\bm{c},q_{l},\nu)$ and the secret key $sk=s$, the plaintext is recovered performing the following computations
$m=\left[\left[c_{0}+c_{1}\cdot s\right]_{q_{l}}\right]_{t}.$
Using the previously established notation, the critical quantity can be rewritten as
$[c_{0}+c_{1}\cdot s]_{q_{l}}=\left[m+t(e\cdot u+e_{1}\cdot s+e_{0})\right]_{q_{l}}=\left[m+t\epsilon\right]_{q_{l}},$
where $t\epsilon$ denotes the error introduced during encryption.
@@
By considering the reduction modulo $t$ of the critical quantity, it is possible to verify that the plaintext is successfully recovered. However, if the error is too large, the value $m+t(e\cdot u+e_{1}\cdot s+e_{0})$ could wrap around the modulus, resulting in an incorrect decryption.
So, decryption is correct only if the error magnitude remains below a certain threshold, as previously emphasized. In addition, the error associated to the ciphertext $\bm{c}$ at level $l$ increases each time the homomorphic circuit is evaluated. Therefore, for constructing a leveled homomorphic encryption scheme, it is crucial to estimate and manage this error properly. To deal with the error effectively, the magnitude of the critical quantity, often referred to as noise, is typically analyzed using its norm. In the context of homomorphic encryption, the infinity and the canonical norm, a definition of which are briefly recalled below [19], assume a central role.
@@
The infinity norm of a polynomial $p\in\mathcal{R}$, of the form $p=p|_{0}+p|_{1}x+...+p|_{\phi(m)-1}x^{\phi(m)-1}$ is defined as
$\|p\|_{\infty}=\max_{0\leq i<\phi(m)}|\hskip 2.84544ptp|_{i}\hskip 2.84544pt|.$
Definition 4
The canonical embedding norm of a polynomial $p\in\mathcal{R}$ is defined as
$\|p\|_{can}=\max_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}|p(\zeta^{j})|,$
where $\zeta$ is a primitive $m$-th root of unity.
Essentially, it corresponds to the infinity norm of the canonical embedding of $p$, denoted as $\sigma(p)$, which is the $\phi(m)$-vector defined as
$\sigma(p)=(p(\zeta^{i}))_{i},$
where $1\leq i<m$ and $\gcd(j,m)=1$.
A relationship between these two norms is given by the following property
$||p||_{\infty}\leq\alpha_{m}||p||_{can}\hskip 28.45274pt\forall p\in\mathcal{R},$
where $\alpha_{m}$ is a constant related to the ring $\mathcal{R}$, which depends exclusively on $m$. Another property, which will be useful in the context of the noise growth analysis, is recalled below [15].
Proposition 2
Let $p,q$ be two polynomials in $\mathcal{R}$. Then,
$||p\cdot q||_{\infty}\hskip 3.41418pt\leq\phi(m)||p||_{\infty}\cdot||q||_{\infty}$
$||p\cdot q||_{can}\leq||p||_{can}\cdot||q||_{can}$
In light of this, to guarantee the correctness of the decryption, the condition on the critical quantity can be expressed as follow
$||\nu||_{\infty}\leq\alpha_{m}||\nu||_{can}<\frac{q_{l}}{2},$
which allows to avoid the error wrapping around the modulus [24]. Naturally, in order to bound the noise, any type of norm could be used. Nevertheless, in the state-of-the-art analysis of noise growth for homomorphic schemes, the canonical norm is typically preferred, due to its properties, as it permits obtaining more accurate estimates. Finally, another concept that is often introduced for the error growth estimation is the noise budget.
Definition 5
Let $(\bm{c},q_{l},\nu)$ be an extended ciphertext. The noise budget associated to $\bm{c}$ is the quantity
$\log_{2}(q_{l})-\log_{2}(||\nu||)-1,$
where |||||$\cdot$||| | ⋅ | | refers to a fixed norm.
The key operations involved in the BGV scheme are briefly summarized in the following.
3.4 Ciphertext Addition and Constant Multiplication
Assume that $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$ are two ciphertexts defined with respect to the same modulus $q_{l}$ and computed encrypting two plaintexts $m$ and $m’$ ’, respectively, using the same key. Their homomorphic sum is defined as
$\mathsf{Add}(\bm{c},\bm{c^{\prime}}):=(c_{0}+c_{0}^{\prime},\;c_{1}+c_{1}^{\prime})\mod q_{l}.$
The resulting ciphertext encrypts the sum $m+m^{\prime}$, and the corresponding noise grows additively. In fact, the resulting critical quantity is defined as
$\nu_{\mathsf{add}}=\nu+\nu^{\prime},$
where $\nu=c_{0}+c_{1}s$ and $\nu^{\prime}=c_{0}^{\prime}+c_{1}^{\prime}s$ denote the critical quantities of $\bm{c}$ and $\bm{c^{\prime}}$, respectively.
Similarly, let $\bm{c}=(c_{0},c_{1})$ be a ciphertext and $k\in R_{t}$ a constant polynomial. The homomorphic multiplication by $k$ is performed as follows
$\mathsf{Mul}_{k}(\bm{c}):=(kc_{0},\;kc_{1})\mod q_{l},$

### v3

@@ Abstract
For an optimal parameter choice, it is crucial to accurately estimate the noise growth, particularly that arising from multiplication, which is the most complex operation. In this work, we propose a novel average-case approach that precisely models noise evolution and guides the selection of initial parameters, improving efficiency while ensuring security. The key innovation of our method lies in accounting for the dependencies among ciphertext errors generated with the same key, and in providing general guidelines for accurate parameter selection that are library-independent.
@@ 1 Introduction
The homomorphic encryption schemes currently in use base their security on the presumed intractability of the Learning with Errors (LWE) problem [36], and its ring variant (RLWE) [33]. Informally, the decisional version of RLWE consists of distinguishing polynomial equations $(a,b=s\cdot a+e)\in\mathcal{R}_{q}\times\mathcal{R}_{q}$, perturbed by small noise $e$ (also called error), from uniform random tuples from $\mathcal{R}_{q}\times\mathcal{R}_{q}$, where $\mathcal{R}_{q}=\mathbb{Z}_{q}[x]/\langle x^{n}+1\rangle$ and $q$ is a positive integer.
@@
Interest in this method, initially applied in the TFHE scheme [13], and subsequently in the CKKS [16, 5], BGV [18, 35] and BFV [4] schemes, grew due to a recognized discrepancy between the estimates based on worst-case technique and experimental data, as highlighted in [17]. The introduction of the average-case approach, as seen in [4, 18], offers a potential resolution to these disparities, indeed, with this method, it is possible to compute a tight probabilistic upper bound.
However, a recent work [24] criticizes current approaches by showing that the noise coefficients may follow a heavy-tailed distribution, potentially invalidating the assumptions underlying average-case bounds. In particular, focusing on BGV without modulus switching, the authors observe a heavy-tailed behavior of the noise. In the following, we confirm this behavior for the same setting, but we further prove that when modulus switching is applied, this is no longer the case. Specifically, under typical parameter choices used by all major libraries, the noise coefficients in BGV with modulus switching follow a Gaussian distribution, and therefore the average-case noise analysis remains valid. Similarly to our demonstration for BGV, a recent work [5] confirms that in CKKS the average-noise analysis remains valid in practice, in line with the observations of [24].
It is worth noting that, the heuristics used for the BGV [35] and CKKS [16] schemes often underestimate the noise growth due to the assumption of the noises independence, leading to imprecise bounds, as also pointed out in [3, 16, 20, 35]. Such underestimates lead to two potential issues: first, the ciphertext is not correctly decrypted with non-negligible probability due to excessive noise and, second, the scheme is exposed to security vulnerabilities, as shown in recent papers [9, 10].
@@
In this paper, we propose the first average-case noise analysis for BGV that does not provide underestimates, taking into account the dependencies introduced by the common secret and public key. We extend the approach of BFV [4], where the authors introduced a correction function $F$ to adjust the product of variances in homomorphic multiplication, by incorporating the dependency effects of the secret key. Similarly, we introduce a correction function which, unlike in [4], is no longer heuristic but derived from formal results presented in our work. Moreover, we observed that in BGV it is necessary to account for additional dependencies introduced by the public key, requiring the correction function to also compensate for these effects. The results obtained in this study suggest that this approach leads to significant improvements in noise analysis.
@@
The paper is structured as follows. Section 2 introduces essential definitions and fundamental properties that are necessary for understanding both our contribution and the functioning of the scheme. Section 3 provides a concise overview of the main features and structure of the BGV scheme. In Section 4, we present our key results concerning the behavior of the error term and its growth under homomorphic operations. Section 5 provides a discussion on why, in the BGV scheme, the error coefficients can be modeled as samples from a discrete Gaussian distribution. Section 6 then demonstrates how the findings from Section 4 and Section 5 can be leveraged to estimate error growth in fixed-operation circuits and to properly select the ciphertext moduli. Section 7 compares our approach with state-of-the-art methods, showing how our parameter selection leads to a significant improvement over those currently adopted in major libraries such as OpenFHE and HElib. Finally, Section 8 concludes the paper and outlines possible directions for future research inspired by our results.
@@ 2 Preliminaries
In this section, we define the general notation and provide the mathematical background that we will use throughout the paper.
2.1 Notation
Let $\mathbb{Z}$ be the ring of integers, and for $d\in\mathbb{Z}_{>0}$ we denote by $\mathbb{Z}_{d}=\nicefrac{{\mathbb{Z}}}{{d\mathbb{Z}}}$. Let $n$ be a power of 2. We define $\mathcal{R}$ as the ring $\mathcal{R}=\mathbb{Z}[x]/\langle x^{n}+1\rangle$ and $\mathcal{R}_{d}=\mathbb{Z}_{d}[x]/\langle x^{n}+1\rangle$.
We use $t$ and $q$ to represent the plaintext and the ciphertext modulus, respectively, and $\mathcal{R}_{t}$ defines the plaintext space, where $t$ is chosen such that $t\equiv 1\mod 2n$. Moreover, to define the ciphertext space, we need to select $L=M+1$ moduli, where $M$ is the multiplicative depth of the circuit. Then, for each level $\ell\in\{0,\dots,L-1\}$, we denote
$q_{\ell}=\prod_{j=0}^{\ell}p_{j},$
as the ciphertext modulus at level L1L-1-$\ell$. Sometimes, the ciphertext modulo $q_{L-1}$ at level 0 will be indicated simply by $q$.
@@ 2.1 Notation
We use lowercase letters such as $a$ for polynomials and bold letters, like $\bm{a}$, for vectors of polynomials. Moreover, we denote by $a|_{i}$ the coefficient of $x^{i}$ of the polynomial $a$. For $a\in\mathcal{R}_{d}$, we let $a]_{d}$ define its centered reduction modulo $d$, with coefficients in [d/2,d/2)[-d/2,d/2). Sometimes, we will also write $a\mod d$. Unless stated otherwise, we assume that the coefficients of the polynomials in $\mathcal{R}_{d}$ are always centered modulo $d$.
@@
Finally, we recall that for $a,b\in\mathcal{R}$ the $i$-th coefficient of their product is given by (see, e.g., [4])
$ab|_{i}=\sum_{j=0}^{n-1}\xi(i,j)a|_{j}b|_{i-j},\quad\mbox{ where }\quad\xi(i,j)=\left\{\begin{array}[]{rl}1&\mbox{ for }i-j\in[0,n)\\ -1&\mbox{ otherwise}.\end{array}\right$. (1)
2.2 Probabilistic distributions
Given a polynomial $a\in\mathcal{R}$ and a probabilistic distribution $\chi$, the notation $a\leftarrow\chi$ is used to indicate that each coefficient of $a$ is randomly and independently sampled according to $\chi$. Some distributions that will be frequently considered are the following:
  • $\mathcal{U}_{q}$ as the uniform distribution over $\mathbb{Z}_{q}$ where the representatives modulo $q$ are taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$;
  • $\mathcal{N}(0,\sigma^{2})$ as the normal distribution, also referred to as the Gaussian distribution, over $\mathbb{R}$, with mean 0 and variance $\sigma^{2}$;
  • $\mathcal{DG}_{q}(\sigma^{2})$ as the discrete Gaussian distribution, which involves sampling a value according to $\mathcal{N}(0,\sigma^{2})$, rounding it to the nearest integer and then reducing it modulo $q$. Moreover, the representative modulo $q$ is taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$.
For the BGV scheme, the notation $\chi_{e}$ and $\chi_{s}$ will be adopted in order to indicate the distribution of the error for a RLWE instance and the secret key coefficients, respectively. Typically, $\chi_{e}$ is a discrete Gaussian distribution with a suitable standard deviation, while for the secret key, the ternary uniform distribution $\mathcal{U}_{3}$ is usually considered. However, in some special cases, where bootstrapping is needed, the choice for the secret key distribution falls on the Hamming Weight distribution where the secret key is sampled uniformly among sparse ternary vectors with a fixed number of non-zero entries.
Moreover, we denote by $V_{a}$ the variance of the coefficients of the polynomial $a$, namely $\mathsf{Var}(a|_{i})$. Given $\gamma\in\mathbb{Z}$, if $a,b\in\mathcal{R}$ are two independent polynomials whose coefficients are independent, identically distributed, and have zero mean, then [17]:
  • $V_{a+b}=V_{a}+V_{b}$
  • $V_{\gamma a}=\gamma^{2}V_{a}$
  • $V_{a\cdot b}=nV_{a}\cdot V_{b}$.
@@ 2.2 Probabilistic distributions
Finally, the values of the variance for some common distributions, which will often be employed in the BGV scheme, are
@@
2.3 Infinity and canonical norms
To conclude this section, we recall the definitions of the infinity and the canonical norm, along with some of their properties.
Definition 1
The infinity norm of a polynomial $a\in\mathcal{R}$ is defined as
$\|a\|_{\infty}=\max_{0\leq i<n}|\hskip 2.84544pta|_{i}\hskip 2.84544pt|.$
If $a\in\mathcal{R}_{q}$ and its coefficients are well-approximated by identically distributed independent Gaussian variables centered at zero, then
$\mathbb{P}\left(\|a\|_{\infty}>T\right)\leq n\left(1-\text{erf}\left(\frac{T}{\sqrt{2V_{a}}}\right)\right),$ (2)
where $\text{erf}(z)$ is the error function, defined as $\text{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\,dt$, see, e.g., [4].
Definition 2
The canonical embedding norm of a polynomial $a\in\mathcal{R}$ is
$\|a\|_{can}=\max_{\begin{subarray}{c}1\leq j<2n\\ \gcd(j,2n)=1\end{subarray}}|a(\zeta^{j})|,$
where $\zeta$ is a primitive $2n$-th root of unity. Essentially, it corresponds to the infinity norm of the canonical embedding of $a$, denoted as $\sigma(a)$.
[…]
We know that if $a\in\mathcal{R}_{q}$ is a random polynomial with coefficient variance $V_{a}$ and $\zeta$ be a primitive $2n^{th}$ root of unity, then the distribution of $a(\zeta)$ is well approximated by a centred Gaussian distribution with variance $nV_{a}$ [20]. This immediately translates into a bound on the canonical norm of $a$:
$a||^{can}<D\sqrt{nV_{a}}\,$, (3)
which holds with probability $(1-e^{-D^{2}/2})^{n}\approx 1-ne^{-D^{2}/2}$. This means that, for a suitable choice of $D$, the bound fails only with negligible probability [20].
@@ 3 The BGV scheme
In this section, we recall the three basic encryption functions of the BGV scheme and the homomorphic operations.
3.1 Basic encryption functions
Key Generation.
The key generation function generates $s\leftarrow\chi_{s}$ , $a\leftarrow\mathcal{U}_{q_{L-1}}$ and $e\leftarrow\chi_{e}$ and outputs the secret key: $\mathsf{sk}=s\in\mathcal{R}_{q_{L-1}}$, and the public key $\mathsf{pk}=(b,a)\equiv(-a\cdot s+te,a)\mod q_{L-1}$.
Encryption.
@@ 3.1 Basic encryption functions
Given the plaintext $m\in\mathcal{R}_{t}$ and the public key $\mathsf{pk}=(b,a)$, the encryption function outputs the ciphertext $\bm{c}\in\mathcal{R}_{q_{L-1}}^{2}$ defined as
$\bm{c}=(c_{0},c_{1})\equiv(b\cdot u+te_{0}+m,a\cdot u+te_{1})\mod q_{L-1},$
where $u,e_{0},e_{1}\in\mathcal{R}_{q_{L-1}}$, with coefficients distributed as $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
Decryption.
@@
Given a ciphertext $\bm{c}\in\mathcal{R}_{q_{\ell}}^{2}$ and the secret key $\mathsf{sk}=s$, the plaintext is recovered performing the following computations $m=\left[\left[c_{0}+c_{1}\cdot s\right]_{q_{\ell}}\right]_{t}$.
We denote by $\nu=[c_{0}+c_{1}\cdot s]_{q_{\ell}}$ the critical quantity corresponding to $\bm{c}$. In particular, for a fresh ciphertext, it can be rewritten as
$\nu_{\mathsf{clean}}=[c_{0}+c_{1}\cdot s]_{q_{L-1}}=[m+t(e\cdot u+e_{1}\cdot s+e_{0})]_{q_{L-1}}=[m+t\epsilon]_{q_{L-1}},$
where $\epsilon$ denotes the error introduced during encryption.
@@
Considering the reduction modulo $t$ of the critical quantity, it is possible to verify that the plaintext is successfully recovered. However, if the error is too large, the value $m+t\epsilon$ could wrap around the modulus, resulting in an incorrect decryption. So, the decryption is correct only if the coefficients of $m+t\epsilon$ remain below a certain threshold.
In addition, the error associated to the ciphertext increases through homomorphic operations [15]. Therefore, it is crucial to estimate the magnitude of the critical quantity (called also noise) which is typically analyzed using its norm.
In light of this, to guarantee the correctness of the decryption, the condition on the critical quantity can be expressed as follows [34]:
[…]
Naturally, in order to bound the noise, any type of norm could be used.
Finally, another concept that is often introduced for the estimation of error growth is the noise budget, which represents the number of bits remaining before wrap-around would occur.
We will use the term extended ciphertext to refer to the tuple $\mathfrak{c}=(\bm{c},q_{\ell},\nu)$ where $\bm{c}$ is the actual ciphertext, $q_{\ell}$ is the ciphertext modulus, and $\nu$ is the associated critical quantity.
@@
[…]
3.2 Homomorphic Operations
Ciphertext Addition.
Let $(\bm{c},q_{\ell},\nu)$ and $(\bm{c}^{\prime},q_{\ell},\nu^{\prime})$ be two extended ciphertexts. Their homomorphic sum is
$\mathsf{Add}(\bm{c},\bm{c^{\prime}}):=(c_{0}+c_{0}^{\prime},\;c_{1}+c_{1}^{\prime})\mod q_{\ell}.$
with resulting critical quantity $\nu_{\mathsf{add}}=\nu+\nu^{\prime}$.
Constant Multiplication.
Let $(\bm{c},q_{\ell},\nu)$ be an extended ciphertext and $k\in\mathcal{R}_{t}$ a fixed polynomial. The homomorphic multiplication by $k$ is
$\mathsf{Mul}_{k}(\bm{c}):=(kc_{0},\;kc_{1})\mod q_{\ell},$